Jennifer, Paul and John are all siblings. Together their ages total 114. John is 5 years younger than Paul and Paul is 2 years younger than Jennifer. Define a variable and then write an equation to find the ages of Jennifer, Paul and John ?
Answer by addingup(3677)

Let John be x, Paul be y, and Jennifer be z
x = y-5
z = y+2
x+y+z = 114 Now substitute per above equalities:
y+y-5+y+2 = 114
3y-3 = 114
3y = 117
y = 39 Paul is 39
x = y-5 = 39-5 = 34 John is 34
z = y+2 = Jennifer is 41
39+34+41 = 114 Correct